# Resurgent Wilson loops in refined topological string

### v1

@@ 1 Introduction
It is now a general agreement that most asymptotic series in physics are in want of a non-perturbative completion. Topological string theory as a subsector of type II superstring theory compactified on Calabi-Yau threefolds has many asymptotic series, which are both mathematically well-defined and more amenable to calculations, and it is therefore a perfect laboratory to explore non-perturbative completions of asymptotic series. In general, the non-perturbative completion would be ambiguous. However, under the assumption that the asymptotic series in topological string is resurgent, the non-perturbative corrections are strongly constrained, and the powerful method of resurgence theory Ecalle can be used to study them.
@@
can be encoded in a trans-series of the form
@@
Furthermore, these trans-series are closedly related to the perturbative series via Stokes transformations, so that much information of the trans-series can be extracted from the perturbative series $\varphi$. For instance, the non-perturbative action $\mathcal{A}_{*}$ that determines the magnitude ${\rm e}^{-\mathcal{A}_{*}/z}$ of the non-perturbative correction is given as singularities of the Borel transform of the asymptotic series, while the coefficients in the non-perturbative corrections, together with the coefficients of the Stokes transformations known as the Stokes constants, can be read off from the large order asymptotics of the perturbative coefficients. These data are sometimes collectively called the resurgent structure.
@@
Recently, these progress have been generalised Alexandrov:2023wdj (see also Alim:2022oll ; Grassi:2022zuk ) to refined free energies in topological string $F(t;\epsilon_{1},\epsilon_{2})$ [^4] The Nekrasov-Shatashvili limit of the refined topological string is special, and the non-perturbative corrections to free energies were studied in Codesido:2017jwp ; CodesidoSanchez:2018vor ; Gu:2022fss . The refined free energy is treated as an asymptotic series in $g_{s}$ denoted as $F(t,\mathsf{b};g_{s})$, where the parameters of Omega background $\epsilon_{1},\epsilon_{2}$ are related to $g_{s}$ via [^5] Our convention differs from that in Alexandrov:2023wdj by $g_{s}\rightarrow{\mathsf{i}}g_{s}$.
@@
In this paper, we would like to study the non-perturbative corrections of another important computable in refined topological string theory, the Wilson loops. They play the role of eigenvalues in the quantum mirror curvse Aganagic:2011mi in the Nekrasov-Shatashvili limit of the refined topological string, and they are encoded in the $qq$-character in the generic Omega background Nekrasov:2015wsu . Topological string on a local Calabi-Yau threefold engineers a 5d $N=1$ SCFT, and when the latter has a gauge theory phase, the Wilson loops are naturally defined. The notion of Wilson loops was extended to topological string on generic local Calabi-Yau threefolds Kim:2021gyj as insertion of additional non-compact 2-cycles, and furthermore, holomorphic anomaly equations for the generalised Wilson loops were written down in Huang:2022hdo ; Wang:2023zcb .
@@
The rest of the paper is structured as follows. We review the resurgent structure of free energy of refined topological string in Section 2, the definition and the calculation of perturbative Wilson loops in Section 3. We present our results of the resurgent structures of Wilson loops in Section 4. We show how to reduce to the unrefined limit and the NS limit in Section 4.2, and we are able to reproduce the results in Gu:2022fss and in particular prove some empirical observations, especially the one that the Stokes constants of NS Wilson loops and those of NS free energies were identical, when the former were not vanishing. Two examples of local $\mathbb{P}^{2}$ and local $\mathbb{P}^{1}\times\mathbb{P}^{1}$ are given in Sections 4.3 and 4.4.
@@ 2.1 Perturbative refined free energy
The coefficients $\mathcal{F}^{(n,g)}(\bm{t})$ are sections of certain line bundles of the moduli space, parametrised by flat coordinates $\bm{t}$.
@@ 3 Perturbative Wilson loops in topological string
However, most of the 5d SCFTs are non-Lagrangian and they do not have a gauge theory phase. Nevertheless, the definition of half-BPS Wilson loops can be generalised through geometric engineering Kim:2021gyj . From the topological string point of view, the Wilson loops arise from the insertion of a collection of non-compact 2-cycles $J=\{\mathsf{C}_{1},\mathsf{C}_{2},\ldots\}$ with infinite volume intersecting with compact 4-cycles in $X$, and the partition function of the topological string on the new threefold $\hat{X}$ with insertion now reads
@@
$Z_{\infty}=Z_{\emptyset}\cdot\exp\sum_{m=1}^{\infty}\frac{1}{m!}F_{m}M^{m},$ (3.11)
@@
$\lim_{N\rightarrow\infty}{N\choose m}F_{m}M_{\mathsf{C}_{1}}^{m}=\lim_{N\rightarrow\infty}\frac{1}{m!}F_{m}(NM_{\mathsf{C}_{1}})^{m}=:\lim_{N\rightarrow\infty}\frac{1}{m!}F_{m}M^{m}.$ (3.12)
@@ 4.1 Solutions to non-perturbative corrections
Finally, the Stokes constant $\mathsf{S}_{\gamma}(\mathsf{b},M)$ for the generating function associated to the non-perturbative action $\mathcal{A}_{\gamma}$ should in principle depend on both $\mathsf{b}$ and $M$. But as the insertion mass $M$ is infinitely heavy while the Stokes constant should be finite, the dependence of the Stokes constant on $M$ should drop out, and we conjecture that
@@ 4.2.2 The NS limit
More generically if the $\mathcal{A}$ is not a local flat coordinate, the non-perturbative refined series in the generic $\ell$-instanton sector in the $\mathsf{b}$-parametrisation are given by (2.28) with the boundary condition (2.27). To derive the NS limit, we require several modifications. We drop the part $\mathcal{F}_{\mathcal{A},1/\mathsf{b}}^{(\ell)}$ associated to the Borel singularities $\ell\mathsf{b}\mathcal{A}$ which will run away to infinity. On the other hand, we use a more generic boundary codition in order to compare with Gu:2022fss ,
$Z_{r,\mathcal{A}}=\exp\left[\sum_{\ell\geq 1}\tau_{\ell}\mathcal{F}_{\mathcal{A},\mathsf{b}}^{(\ell)}\right]=1+\sum_{n>0}C_{n}(\tau)\exp\left(-\frac{n\mathcal{A}}{\mathsf{b}g_{s}}\right).$ (4.37)
The generic non-perturbative component of the refined partition function is
@@
As the non-perturbative corrections to the NS free energy $F_{\text{NS}}(\bm{t};\hbar)$ are functions of the free energy itself, the non-perturbative corrections to the generating series (4.46) is obtained by the same expression but with the NS free energy replaced by the NS generating series. The non-perturbative corrections to individual NS BPS sectors are extracted as coefficients of $M^{m}/m!$. For instance, from (4.2.2), we immediately get
@@
Two important features of the resurgent structures of the Wilson loop veves were discovered in Gu:2022fss : the Borel singularities are integral periods which are not local flat coordinates, and that the Stokes constants, if they are not vanishing, are the same as those of the NS free energy. And they were presented as independent results from those of the NS free energies. It is clear now that they are directly related, as both the NS Wilson loop and the NS free energy are parts of the same generating series, which enjoys the uniform resurgent structure.
@@ 4.3 Example: local $\mathbb{P}^{2}$
We consider the local $\mathbb{P}^{2}$, i.e. the total space of the canonical bundle of $\mathbb{P}^{2}$, in this section. Local $\mathbb{P}^{2}$ is a basic Calabi-Yau manifold, but with rich geometric structure, and it has been discussed in great detail in the literature. We follow the convention in Haghighat:2008gw ; Couso-Santamaria2015 where its moduli space is parametrised by the global complex coordinate $z$ such that the large radius singularity, the conifold singularity, and the orbifold singularity are located respectively at $z=0$, $z=-1/27$ and $z=\infty$.
@@
Figure 4.1: Borel singularities of refined Wilson loop BPS sectors $\mathcal{F}[1](\mathsf{b};g_{s})$ for local $\mathbb{P}^{2}$ with $\mathsf{b}=2$ up to $g=50$ in the large radius frame, respectively (a) near the large radius point $z=0$ and (b) near the conifold point $z=-1/27$. The red dots are the approximate singularities from numerical calculations, and the black dots are the indicated branch points, located at (a) $\mathsf{b}^{-1}\mathcal{A}_{(-3,0,0)_{\text{LR}}},\mathsf{b}^{-1}\mathcal{A}_{(-3,1,0)_{\text{LR}}},$ and (b) $\mathsf{b}^{-1}\mathcal{A}_{(-3,0,0)_{\text{LR}}}$ respectively.
@@
Figure 4.2: Borel singularities of refined Wilson loop BPS sectors $\mathcal{F}[2](\mathsf{b};g_{s})$ for local $\mathbb{P}^{2}$ with $\mathsf{b}=2$ up to $g=50$ in the large radius frame, respectively (a) near the large radius point $z=0$ and (b) near the conifold point $z=-1/27$. The red dots are the approximate singularities from numerical calculations, and the black dots are the indicated branch points, located at (a) $\mathsf{b}^{-1}\mathcal{A}_{(-3,0,0)_{\text{LR}}},\mathsf{b}^{-1}\mathcal{A}_{(-3,1,0)_{\text{LR}}},$ and (b) $\mathsf{b}^{-1}\mathcal{A}_{(-3,0,0)_{\text{LR}}}$ respectively.
@@
Figure 4.3: Borel singularities of refined free energies $\mathcal{F}[0](\mathsf{b};g_{s})$ for local $\mathbb{P}^{2}$ with $\mathsf{b}=2$ up to $g=50$ in the large radius frame, respectively (a) near the large radius point $z=0$ and (b) near the conifold point $z=-1/27$. The red dots are the approximate singularities from numerical calculations, and the black dots are the indicated branch points, located at (a) $\mathsf{b}^{-1}\mathcal{A}_{(-3,0,0)_{\text{LR}}}$, $\mathsf{b}^{-1}\mathcal{A}_{(0,1,n)_{\text{LR}}}(n=0,1,2)$ and (b) $\mathsf{b}^{-1}\mathcal{A}_{(-3,0,0)_{\text{LR}}}$ respectively.
@@
Figure 4.4: Borel singularities of refined Wilson loop BPS sectors $\mathcal{F}[1](\mathsf{b};g_{s})$ for local $\mathbb{P}^{2}$ with $\mathsf{b}=2$ up to $g=50$ in the conifold frame, respectively (a) near the large radius point with $z<0$, (c) near the conifold point, and (c) with $z>0$. The red dots are the approximate singularities from numerical calculations, and the black dots are the indicated branch points, located at (a) $\mathsf{b}^{-1}\mathcal{A}_{(-3,0,n)_{\text{c}}}(n=0,1,2)$, (b) $\mathsf{b}^{-1}\mathcal{A}_{(-3,1,0)_{\text{c}}}$ and (c) $\mathsf{b}^{-1}\mathcal{A}_{(-3,0,0)_{\text{c}}}$ (right), $\mathsf{b}^{-1}\mathcal{A}_{(-3,1,0)_{\text{c}}}$ (below) respectively.
@@
Figure 4.5: Borel singularities of refined Wilson loop BPS sectors $\mathcal{F}[2](\mathsf{b};g_{s})$ for local $\mathbb{P}^{2}$ with $\mathsf{b}=2$ up to $g=50$ in the conifold frame, respectively (a) near the large radius point with $z<0$, (b) near the conifold point $z=-1/27$, and (c) with $z>0$. The red dots are the approximate singularities from numerical calculations, and the black dots are the indicated branch points, located at (a) $\mathsf{b}^{-1}\mathcal{A}_{(-3,0,n)_{\text{c}}}(n=0,1,2)$, (c) $\mathsf{b}^{-1}\mathcal{A}_{(-3,1,0)_{\text{c}}}$, and (c) $\mathsf{b}^{-1}\mathcal{A}_{(-3,0,0)_{\text{c}}}$ (right), $\mathsf{b}^{-1}\mathcal{A}_{(-3,1,0)_{\text{c}}}$ (below) respectively.
@@
as shown in Figs. 4.4 (c), 4.5 (c). We find yet again that none of the Borel singularities coincide with the flat coordinate up to a constant, i.e. the coefficient $c^{\prime}$ in $\gamma_{\text{c}}$ does not vanish.
@@
Figure 4.7: Comparison for local $\mathbb{P}^{2}$ for local $\mathbb{P}^{2}$ between numerical results (red dots) of $\frac{1}{\pi{\mathsf{i}}}\mathsf{S}\cdot\mu_{0,1,2}$ extracted from the large order asymptotics of $\mathcal{F}_{g}[2]$ up to $g=50$ in the large radius frame at $\mathsf{b}=2$ with error bars (vertical bars, virtually invisible) and trans-series solutions from HAE (solid line). Richardson transformation of degree 10 is used to improve the numerics. The horizontal axis is modulus $z$.
@@
We consider two different cases. The first is the BPS sectors near the conifold point in the large radius frame. The dominant Borel singularities are the pair of $\gamma_{\text{LR}}=\pm(-3,0,0)$ as shown in Figs. 4.1 (b), 4.2 (b). The large order asymptotics formula (4.64) (formula (4.66) in the unrefined limit) can be used to extract the non-perturbative coefficients $\mu_{0},\mu_{1},\mu_{2},\ldots$, and we compare these numerical results with our prediction from Sections 4.1 and 4.2.1 in Figs. 4.6, 4.7 for generic $\mathsf{b}$, and in Fig. 4.23 in the unrefined limit. The numerical results and the theoretical predictions match perfectly, as long as we choose the Stokes constants
@@
Similarly, we consider the BPS sectors at $z>0$ in the conifold frame. Depending on the actual value of $z$, the two pairs of Borel singularities $\gamma_{\text{c}}=\pm(-3,1,0)$ and $\gamma_{\text{c}}=\pm(-3,0,0)$ compete in dominance, as shown in Fig. 4.4 (c), 4.5 (c). If $z\gtrsim 4\times 10^{-6}$, the pair of $\gamma_{\text{c}}=\pm(-3,1,0)_{\text{c}}$ is dominant (closer to the origin). The comparison between the numerical results of $\mu_{0},\mu_{1},\mu_{2}$ from large order asymptotics and theoretical predictions are plotted in Figs. 4.10, 4.11. Here we have chosen the Stokes constants
@@
Finally, if $0<z\lesssim 4\times 10^{-6}$, the pair of $\gamma_{\text{c}}=\pm(-3,0,0)_{\text{c}}$ is dominant, and the plots for $\mu_{0},\mu_{1},\mu_{2}$ are given in Figs. 4.12, 4.13. Here we have chosen the Stokes constants

### v2

@@ 1 Introduction
It is now a general agreement that most asymptotic series in physics are in want of a non-perturbative completion. Topological string theory as a subsector of type II superstring theory compactified on Calabi-Yau threefolds has many asymptotic series, such as free energies and Wilson loop vevs, which are both mathematically well-defined and more amenable to calculations, and it is therefore a perfect laboratory to explore non-perturbative completions of asymptotic series. In general, the non-perturbative completion would be ambiguous. However, under the assumption that the asymptotic series in topological string is resurgent, the non-perturbative corrections are strongly constrained, and the powerful method of resurgence theory Ecalle can be used to study them.
@@
can be encoded in the form
@@
Furthermore, these non-perturbative corrections are closedly related to the perturbative series via Stokes transformations, so that much information of the non-perturbative corrections can be extracted from the perturbative series $\varphi$. For instance, the non-perturbative action $\mathcal{A}_{*}$ that determines the magnitude ${\rm e}^{-\mathcal{A}_{*}/z}$ of the non-perturbative correction is given as singularities of the Borel transform of the asymptotic series, while the coefficients in the non-perturbative corrections, together with the coefficients of the Stokes transformations known as the Stokes constants, can be read off from the large order asymptotics of the perturbative coefficients. These data are sometimes collectively called the resurgent structure.
@@
Recently, these progress have been generalized Alexandrov:2023wdj (see also Alim:2022oll ; Grassi:2022zuk ) to refined free energies in topological string $F(t;\epsilon_{1},\epsilon_{2})$ [^4] The Nekrasov-Shatashvili limit of the refined topological string is special, and the non-perturbative corrections to free energies were studied in Codesido:2017jwp ; CodesidoSanchez:2018vor ; Gu:2022fss . The refined free energy is treated as an asymptotic series in $g_{s}$ denoted as $F(t,\mathsf{b};g_{s})$, where the parameters of Omega background $\epsilon_{1},\epsilon_{2}$ are related to $g_{s}$ via [^5] Our convention differs from that in Alexandrov:2023wdj by $g_{s}\rightarrow{\mathsf{i}}g_{s}$.
@@
In this paper, we would like to study the non-perturbative corrections of another important computable in refined topological string theory, the Wilson loops. They play the role of eigenvalues in the quantum mirror curves Aganagic:2011mi in the Nekrasov-Shatashvili limit of the refined topological string, and they are encoded in the $qq$-character in the generic Omega background Nekrasov:2015wsu . Topological string on a local Calabi-Yau threefold engineers a 5d $N=1$ SCFT, and when the latter has a gauge theory phase, the Wilson loops are naturally defined. The notion of Wilson loops was extended to topological string on generic local Calabi-Yau threefolds Kim:2021gyj as insertion of additional non-compact 2-cycles, and furthermore, holomorphic anomaly equations for the generalized Wilson loops were written down in Huang:2022hdo ; Wang:2023zcb .
@@
The rest of the paper is structured as follows. We review the resurgent structure of free energy of refined topological string in Section 2, the definition and the calculation of perturbative Wilson loops in Section 3. We present our results of the resurgent structures of Wilson loops in Section 4. We show how to reduce to the unrefined limit and the NS limit in Section 4.2, and we are able to reproduce the results in Gu:2022fss and in particular prove some empirical observations, especially the one that the Stokes constants of NS Wilson loops and those of NS free energies were identical, when the former were not vanishing. Two examples of local $\mathbb{P}^{2}$ and local $\mathbb{P}^{1}\times\mathbb{P}^{1}$ are given in Section 5.
@@ 2.1 Perturbative refined free energy
The coefficients $\mathcal{F}^{(n,g)}(\bm{t})$ are sections of certain line bundles of the moduli space, parametrized by flat coordinates $\bm{t}$.
@@ 3 Perturbative Wilson loops in topological string
However, most of the 5d SCFTs are non-Lagrangian and they do not have a gauge theory phase. Nevertheless, the definition of half-BPS Wilson loops can be generalized through geometric engineering Kim:2021gyj . From the topological string point of view, the Wilson loops arise from the insertion of a collection of non-compact 2-cycles $J=\{\mathsf{C}_{1},\mathsf{C}_{2},\ldots\}$ with infinite volume intersecting with compact 4-cycles in $X$, and the partition function of the topological string on the new threefold $\hat{X}$ with insertion now reads
@@
$Z_{\infty}=Z_{\emptyset}\cdot\exp\sum_{m=1}^{\infty}\frac{1}{m!}F[m]M^{m},$ (3.11)
@@
$\lim_{N\rightarrow\infty}{N\choose m}F[m]M_{\mathsf{C}_{1}}^{m}=\lim_{N\rightarrow\infty}\frac{1}{m!}F[m](NM_{\mathsf{C}_{1}})^{m}=:\lim_{N\rightarrow\infty}\frac{1}{m!}F[m]M^{m}.$ (3.12)
@@ 4.1 Solutions to non-perturbative corrections
Finally, the Stokes constant $\mathsf{S}_{\gamma}(\mathsf{b},M)$ for the generating function associated to the non-perturbative action $\mathcal{A}_{\gamma}$ should in principle depend on both $\mathsf{b}$ and $M$. But as the insertion mass $M$ is infinitely heavy while the Stokes constant should be a finite number, the dependence of the Stokes constant on $M$ should drop out. In fact, the Stokes constants of the topological string free energies are generally conjectured to be constants independent of the Kähler moduli of the Calabi-Yau threefold Gu:2021ize ; Gu:2022sqc , the Stokes constants for the generating series, which is the topological string free energy of the new threefold $\hat{X}$, should also not depend on the Kähler moduli, $M$ included. We thus conjecture that
@@ 4.2.2 The NS limit
More generically if the $\mathcal{A}$ is not a local flat coordinate, the non-perturbative refined series in the generic $\ell$-instanton sector in the $\mathsf{b}$-parametrisation are given by (2.28) with the boundary condition (2.27). To derive the NS limit, we require several modifications. We drop the part $\mathcal{F}_{\mathcal{A},1/\mathsf{b}}^{(\ell)}$ associated to the Borel singularities $\ell\mathsf{b}\mathcal{A}$ which will run away to infinity. On the other hand, we use a more generic boundary condition in order to compare with Gu:2022fss ,
$Z_{r,\mathcal{A}}=\exp\left[\sum_{\ell\geq 1}\tau_{\ell}\mathcal{F}_{\mathcal{A},\mathsf{b}}^{(\ell)}\right]=1+\sum_{n>0}C_{n}(\tau)\exp\left(-\frac{n\mathcal{A}}{\mathsf{b}g_{s}}\right),$ (4.37)
where $\tau_{\ell}$ are constants that parametrize the boundary condition. The generic non-perturbative component of the refined partition function is
@@
As the non-perturbative corrections to the NS free energy $F_{\text{NS}}(\bm{t};\hbar)$ are functions of the free energy itself, the non-perturbative corrections to the generating series (4.46) is obtained by the same expression but with the NS free energy replaced by the NS generating series. The non-perturbative corrections to individual NS BPS sectors are extracted as coefficients of $M^{m}/m!$. For instance, from (4.41), we immediately get
@@
Two important features of the resurgent structures of the Wilson loop vevs were discovered in Gu:2022fss : the Borel singularities are integral periods which are not local flat coordinates, and that the Stokes constants, if they are not vanishing, are the same as those of the NS free energy. And they were presented as independent results from those of the NS free energies. It is clear now that they are directly related, as both the NS Wilson loop and the NS free energy are parts of the same generating series, which enjoys the uniform resurgent structure.
5 Examples
@@ 5.1 Example: local $\mathbb{P}^{2}$
We consider the local $\mathbb{P}^{2}$, i.e. the total space of the canonical bundle of $\mathbb{P}^{2}$, in this section. Local $\mathbb{P}^{2}$ is a basic Calabi-Yau manifold, but with rich geometric structure, and it has been discussed in great detail in the literature. We follow the convention in Haghighat:2008gw ; Couso-Santamaria2015 where its moduli space is parametrized by the global complex coordinate $z$ such that the large radius singularity, the conifold singularity, and the orbifold singularity are located respectively at $z=0$, $z=-1/27$ and $z=\infty$.
@@
Figure 5.1: Borel singularities of refined Wilson loop BPS sectors $\mathcal{F}[1](\mathsf{b};g_{s})$ for local $\mathbb{P}^{2}$ with $\mathsf{b}=2$ up to $g=50$ in the large radius frame, respectively (a) near the large radius point $z=0$ and (b) near the conifold point $z=-1/27$. The red dots are approximate singularities from numerical calculations, which would accumulate to branch cuts. The branch points (black dots) on the imaginary axis are $\mathsf{b}^{-1}\mathcal{A}_{\pm(-3,0,0)_{\text{LR}}}$, and those in the four quadrants are $\mathsf{b}^{-1}\mathcal{A}_{\pm(-3,1,0)_{\text{LR}}}$ and $\mathsf{b}^{-1}\mathcal{A}_{\pm(-3,-1,-1)_{\text{LR}}}$.
@@
Figure 5.2: Borel singularities of refined Wilson loop BPS sectors $\mathcal{F}[2](\mathsf{b};g_{s})$ for local $\mathbb{P}^{2}$ with $\mathsf{b}=2$ up to $g=50$ in the large radius frame, respectively (a) near the large radius point $z=0$ and (b) near the conifold point $z=-1/27$. The red dots are approximate singularities from numerical calculations, which would accumulate to branch cuts. The branch points (black dots) are the same as in Fig. 5.1.
@@
Figure 5.3: Borel singularities of refined free energies $\mathcal{F}[0](\mathsf{b};g_{s})$ for local $\mathbb{P}^{2}$ with $\mathsf{b}=2$ up to $g=50$ in the large radius frame, respectively (a) near the large radius point $z=0$ and (b) near the conifold point $z=-1/27$. The red dots are approximate singularities from numerical calculations. The branch points (black dots) on the imaginary axis are $\mathsf{b}^{-1}\mathcal{A}_{\pm(-3,0,0)_{\text{LR}}}$, and those away from the imaginary axis are $\mathsf{b}^{-1}\mathcal{A}_{\pm(0,1,n)_{\text{LR}}}(n=-1,0,1,2)$.
@@
Figure 5.4: Borel singularities of refined Wilson loop BPS sectors $\mathcal{F}[1](\mathsf{b};g_{s})$ for local $\mathbb{P}^{2}$ with $\mathsf{b}=2$ up to $g=50$ in the conifold frame, respectively (a) near the large radius point with $z<0$, (c) near the conifold point, and (c) with $z>0$. The red dots are approximate singularities from numerical calculations. The branch points (black dots) are (a) $\mathsf{b}^{-1}\mathcal{A}_{\pm(-3,0,n)_{\text{c}}}(n=-1,0,1,2)$, (b) $\mathsf{b}^{-1}\mathcal{A}_{\pm(-3,1,0)_{\text{c}}}$, $\mathsf{b}^{-1}\mathcal{A}_{\pm(-3,-1,1)_{\text{c}}}$ and (c) $\mathsf{b}^{-1}\mathcal{A}_{\pm(-3,0,0)_{\text{c}}}$ (horizontal), $\mathsf{b}^{-1}\mathcal{A}_{\pm(-3,-1,0)_{\text{c}}}$ respectively.
@@
Figure 5.5: Borel singularities of refined Wilson loop BPS sectors $\mathcal{F}[2](\mathsf{b};g_{s})$ for local $\mathbb{P}^{2}$ with $\mathsf{b}=2$ up to $g=50$ in the conifold frame, respectively (a) near the large radius point with $z<0$, (b) near the conifold point $z=-1/27$, and (c) with $z>0$. The red dots are approximate singularities from numerical calculations. The branch points (black dots) are the same as in Fig. 5.4.
@@
as shown in Figs. 5.4 (c), 5.5 (c). We find yet again that none of the Borel singularities coincide with the flat coordinate up to a constant, i.e. the coefficient $c^{\prime}$ in $\gamma_{\text{c}}$ does not vanish.
@@
We consider two different cases. The first is the BPS sectors near the conifold point in the large radius frame. The dominant Borel singularities are the pair of $\gamma_{\text{LR}}=\pm(-3,0,0)$ as shown in Figs. 5.1 (b), 5.2 (b). The large order asymptotics formula (5.13) (formula (5.15) in the unrefined limit) can be used to extract the non-perturbative coefficients $\mu_{0},\mu_{1},\mu_{2},\ldots$, and we compare these numerical results with our prediction from Sections 4.1 and 4.2.1 in Figs. 5.6, 5.7 for generic $\mathsf{b}$, and in Figs. 5.8, 5.9 in the unrefined limit. The numerical results and the theoretical predictions match perfectly, as long as we choose the Stokes constants
@@
Similarly, we consider the BPS sectors at $z>0$ in the conifold frame. Depending on the actual value of $z$, the two pairs of Borel singularities $\gamma_{\text{c}}=\pm(-3,1,0)$ and $\gamma_{\text{c}}=\pm(-3,0,0)$ compete in dominance, as shown in Fig. 5.4 (c), 5.5 (c). If $z\gtrsim 4\times 10^{-6}$, the pair of $\gamma_{\text{c}}=\pm(-3,1,0)_{\text{c}}$ is dominant (closer to the origin). The comparison between the numerical results of $\mu_{0},\mu_{1},\mu_{2}$ from large order asymptotics and theoretical predictions are plotted in Figs. 5.10, 5.11. Here we have chosen the Stokes constants
@@
Finally, if $0<z\lesssim 4\times 10^{-6}$, the pair of $\gamma_{\text{c}}=\pm(-3,0,0)_{\text{c}}$ is dominant, and the plots for $\mu_{0},\mu_{1},\mu_{2}$ are given in Figs. 5.12, 5.13 (in the unrefined limit). Here we have chosen the Stokes constants